You have agreed to make investment in your friends agricultural farm. This would require an amount of $20,000 as initial investment on your part. Your friend promises you revenue (before expenses) of $3600 per year the first year and thereafter the revenue increases by $200 per year. Your share of the estimated annual expenses is $1000. You are planning to invest for six years. Your friend has made the commitment to buyout your share of the business at that time for $24000. You have decided to set a personal MARR of 15% per year. Judge the profitability of the investment project by using Future Worth (FW) method.


1
Expert's answer
2021-11-25T10:24:07-0500

Solution:

Using NPV:

NPV = "(\sum\frac{ CF}{(1+r)^{n} } ) - Initial \;investment"

CF = Net cash flows

r = Interest rate = 15%

n = Year

Derive net cash flows: Revenues - Profits

Year 1 = 3,600 – 1,000 = 2,600

Year 2 = 3,800 – 1,000 = 2,800

Year 3 = 4,000 – 1,000 = 3,000

Year 4 = 4,200 – 1,000 = 3,200

Year 5 = 4,400 – 1,000 = 3,400

Year 6 = 4,600 – 1,000 = 3,600

 

NPV = "(\frac{ 2,600}{(1+0.15)^{1} } + \frac{ 2,800}{(1+0.15)^{2} } + \frac{ 3,000}{(1+0.15)^{3} } + \frac{ 3,200}{(1+0.15)^{4} } + \frac{ 3,400}{(1+0.15)^{5} } + \frac{ 3,600}{(1+0.15)^{6} }) - 20,000"


NPV = (2260.87 + 2117.20 + 1972.55 + 1829.61 + 1690.40 + 1556.38) – 20,000

NPV = 11,427.01 – 20,000 = (8,572.99)


NPV = (8,572.99)

 

The project NPV is negative, which means that the discounted present value of all future cash flows related to the project will be negative, and therefore unattractive.

Therefore, you should accept your current friend's offer of $24,000.